# Reclaiming “Open AI" - AI Model Serving Can Be Open Access, Yet Monetizable and Loyal

### v3

@@ B.2 Coalition Attack
An adversarial host who has legitimately acquired multiple OMLized models can launch a notorious attack known as coalition attacks, where multiple OMLized models are used to evade fingerprint detection. One such attack is studied in [21] where common model merging techniques including [109, 41, 113, 116] are used against instructional fingerprinting [111] and watermarking [54]. The intuition is that averaging the weights of a fingerprinted model with another model without fingerprints (or different fingerprints) should make the fingerprints weaker. In the promising preliminary results of [21], the fingerprinting techniques of [111] demonstrated robustness against such attacks; fingerprints persisted through all model merging that preserve utility. However, this is a weak attack and can be significantly strengthened. Note that one implication of this robustness of model merging is that it can be used for trust-free OML as we discuss in Section B.5.1. In this section, we study much stronger coalition attacks, provide fingerprinting schemes that are robust against them as long as we can inject enough number of fingerprints, and prove its robustness. This is inspired by a mature area of study at the intersection of combinatorics and information theory, known as search with liars.
@@
Strong coalition attacks. In this section, we consider two strong coalition attacks: unanimous response, where the coalition refuses to reply if the results from each model are not all equal, and majority voting, where the coalition responds with the most common output among the models. Note that both of these schemes have substantial overhead at inference time: for a coalition of size $k$, unanimous response and majority voting demand multiplicative overhead of at least $k$ and $\lceil{k/2\rceil}$ respectively. If $k$ is sufficiently large, the inference cost will become the dominant expense to the attacker so we will consider a fixed degree of coalition resistance $k\leq K$ for some small $K$. Note that these are stronger coalition attacks than the simple model merging studied in [21], which simply merges the weights of the $k$ models; even when each model has distinct fingerprints, model merging attack has been demonstrated to fail. The standard fingerprint schemes are robust against model merging attacks as we show in Section B.5.1. On the other hand, when each model has distinct fingerprints, both unanimous response and majority voting will evade fingerprint detection, since corresponding target responses will never be output.
@@
To address these stronger coalition attacks of unanimous response and majority voting, we design a novel fingerprinting scheme. This is inspired by the literature on search with liars, and we show that, with enough fingerprints, we can provably identify the models participating in the coalition attacks. The main idea is to add each fingerprint to multiple OMLized models in a carefully designed manner, such that we can iteratively narrow down the candidate set of deployed OMLized models that contains all the models in the coalition of interest. Precisely, let the total number of possible deployed OMLized models be $N$ and the maximum coalition size is $K$ (or $12K-$ 1 in the case of majority voting).
Proposition B.1.
@@
There exists a randomized fingerprinting scheme for a universe of $N$ models which can identify a unanimous response coalition of size $K$ (or a majority voting coalition of size $12K-$ 1) using
@@
Proof of Proposition B.1.
The scheme proceeds with leave-one-out fingerprinting for partitioning of the models as follows: In each round, we assume the candidate models have been split into $K+1$ disjoint partitions $P_{1},\ldots,P_{K+1}$ such that $[N]=P_{1}\sqcup\cdots\sqcup P_{K+1}$. Then, for each partition $P_{i}$, we inject one fingerprint $F_{i}$ into each model in the complement $[N]\setminus P_{i}$. When testing for the fingerprint, we check for all $K+1$ possible fingerprints $F_{i}$. This guarantees that there will be a fingerprint $F_{i^{*}}$ which spans the coalition (or the acting majority in the case of majority voting), since the no more than $K$ models that determined the coalition’s output can span at most $K$ distinct partitions. Once we have identified $F_{i^{*}}$, we can eliminate the partition $P_{i^{*}}$ from the candidate set. Our goal will be to recursively apply this procedure until the exact coalition has been identified.
@@
To construct the partitions for all rounds ahead of time, we randomly sample $R$ groups of evenly sized partitions $\{P^{(1)}_{i}\}_{i=1}^{K+1},\ldots,\{P^{(R)}_{i}\}_{i=1}^{K+1}$ uniformly from the space of such partitions (thus, all partitions have size $N/(K+1)$). Although the partitions may not remain evenly sized after the candidate set has been narrowed, we will show that we are still able to make progress in each round. Let $C$ denote the candidate set. Then for any choice of $r$ and $i$, the size of $C\cap P^{(r)}_{i}$ is distributed as $\operatorname{Hypergometric}(N,N/(K+1),\lvert C\rvert)$. Then, by a standard Hypergeometric tail bound, we know that
$\operatorname{P}\left(\left\lvert C\cap P^{(r)}_{i}\right\rvert\leq\left(\frac{1}{K+1}-\zeta\right)\lvert C\rvert\right)\;\;\leq\;\;\exp\left(-2\zeta^{2}\lvert C\rvert\right).$
@@
To shrink, $C$ from size $N$ to $N_{0}$, it is sufficient to have $R_{0}=\log(\frac{N}{N_{0}})/\log(1+\frac{1}{2k+1})=O(K\log N)$ successful rounds. By a binomial tail bound, $O(R_{0}\log(1/\delta))$ rounds are sufficient to guarantee $R_{0}$ successes with probability at least $21-\delta/$ 2. Now, considering the regime where $C$ is shrinking from size $N_{0}$ to 0 (at worst), we note that
@@
In this regime, we define a round a successful if the candidate set shrinks by at least 1. The only way a round can fail is when all partitions that do not contain any coalition members (of which there must be at least one) do not intersect with $C$. From the above, we see that the round must succeed with probability at least $\frac{1}{K+1}$. Now, to successfully identify the coalition, $N_{0}$ successful rounds suffice (we will terminate early once the coalition is identified). By a binomial tail bound, $O(K\cdot N_{0}\log(1/\delta))$ rounds are sufficient to guarantee $N_{0}$ successes with probability at least $21-\delta/$ 2. Combining the rounds from both regimes, we see that $R=O\big{(}(K\log N+K^{3}\log K)\log(1/\delta)\big{)}$ ensures overall success with probability at least $1-\delta$. Finally, recall that each round uses $O(K)$ fingerprints. ∎
Worst-case coalition attacks. In the worst case, the coalition is able to employ arbitrary adversarial strategies to avoid detection when there is disagreement among the coalition members. This is significantly more challenging as the adaptive detection algorithm of Proposition B.1 does not guarantee accurate detection anymore. In general, this problem can be formulated as search with lies [52, 78, 53]. In particular, it follows from [53] that there is no fingerprinting procedure that can deterministically guarantee the identification of the coalition, even when assigning unique fingerprints to all possible subsets of models. (Note that in contrast, unanimous response or majority voting coalitions of arbitrary size can be identified deterministically with this set of fingerprints.) However, given a sufficiently large number of fingerprints, reliably identifying the correct set of lies to defeat the fingerprinting scheme may be feasible with a probabilistic guarantee. We demonstrate this in the following proposition.
Proposition B.2.
@@
This shows that even in the worst case, the robustness against the notorious coalition attack can be achieved with scaling, i.e., as long as we have enough fingerprints. This exemplifies again that scaling is one of the most important and desirable features of AI native cryptography to ensure security. Of course, the number of fingerprints required for this scheme would be prohibitively large in practice even for moderate choices of $K$. Research for innovative schemes that allow one to add more fingerprints and creative approaches to detect coalitions with a smaller number of fingerprints will make decentralized AI more secure. At the same time, we believe this result can be improved with a robust version of an adaptive algorithm similar to the one in Proposition B.1. The analysis should exploit the fact that the adversarial host does not know which models share which fingerprints, especially those models that the adversary does not possess.
Proof of Proposition B.2.
@@
Noting that $K\leq\sqrt{N/2}$ and choosing $M=O(K\log(N/\delta))$ completes the proof. ∎
B.3 Sentient Protocol under Decentralized and Untrusted Provers
In OML 1.0, we say a protocol is secure if a host who does not acquire signed permission strings when using an OMLized model can be detected with high probability. Ideally, we want a protocol that is secure without relying on trusted provers. Given a pool of decentralized provers, we demonstrate that the Sentient protocol is secure as long as at least one of the provers is honest and the fingerprint responses are kept secret.
Threat model. Consider the scenario of Section B.1.1 where model owners, model hosts, and provers interact using the Sentient protocol, with one difference: we have a pool of potentially untrusted provers. Concretely, under the threat model of Section B.1.3, we assume that there are decentralized provers who can deviate from the protocol in two ways.
@@ B.3 Sentient Protocol under Decentralized and Untrusted Provers
Secondly, an adversarial prover can fabricate a proof of usage to frame an honest host. When an adversarial prover reports a fabricated key-response pair, $(\tilde{q},M.{\rm oml}(\tilde{q}))$, without querying the host, the previous Sentient protocol that trusts provers has no way of telling whether the prover is lying or the host has not acquired the signed permission.
Security analysis under decentralized and untrusted provers. To address these two attacks, we assume that the Sentient protocol ensures that ($i$) there is at least on honest prover in the pool, ($ii$) the provers have access to only the fingerprint keys, $\{\tilde{q}\}$, and not the target responses, $\{M.{\rm oml}(\tilde{q})\}$, and ($iii$) each prover only has access to a disjoint subset of the fingerprint keys.
@@
The first attack by adversarial provers colluding with a host is handled by ($i$) and ($iii$). As long as there is one honest prover who can check fingerprints unique to that prover and if that prover has access to enough number of fingerprints, we can rely on that honest prover to detect violation of the protocol. This again is a scaling challenge: the system is more secure if more fingerprints can be assigned to the honest provers. As long as we have enough fingerprints assigned to the honest provers, robustness of our fingerprints to input perturbation (Section B.1.3) and fine-tuning (Section B.1.4) will still hold.
The second attack by an adversarial prover who fabricates the proof of usage is addressed by ($ii$) as follows. The verification step in Figure 10 is robust against fabricating a proof of usage as long as the prover does not know the target response to the key, $\tilde{q}$, and the target response chosen for the fingerprint is difficult to guess (with low enough probability of successfully guessing it). This ensures that it is nearly impossible for a prover to fabricate the fingerprint response paired with $\tilde{q}$ without actually running inference on the host’s model, and such an unmatched proof of usage, $(\tilde{q},\tilde{r})$ will be rejected by the verifier in Step 3 of Figure 10.
@@
For coalition attacks, our schemes in Section B.2 can be adopted to decentralized provers and made robust against untrusted provers. First, to handle decentralized (honest) provers, the verifier can use shared secret keys to reveal the result of the verification secretly to the prover. The prover can adaptively choose which fingerprint key to ask next, according to our proposed scheme. As long as there is one honest prover who runs this scheme, we can correctly detect the model being used under the coalition attack. Note that an adversarial prover can only cause false negatives, i.e., turn a positive proof of usage into a negative proof. The non-adaptive fingerprinting scheme of Proposition B.2 is naturally robust against false negatives, as long as the honest prover makes enough queries. The adaptive fingerprinting scheme of Proposition B.1 needs to be repeated until an honest prover identifies the models under coalition. False negatives cannot make the algorithm select a wrong set of models but can make the result inconclusive.
B.4 Achieving Loyalty in OML 1.0
@@
The Sentient protocol for OML 1.0 introduced in this paper addresses Openness and Monetization, but not Loyalty. One of the most important applications of loyalty is the alignment of LLMs to human safety preferences. Recent advances in hardening the models to be robustly aligned against fine-tuning and jail-breaking attacks can shed light on how to achieve Loyalty on top of OML 1.0.
@@
B.5 Discussion
B.5.1 Trust-free OML 1.0
Ideally, we want OML to not rely on the trust of any party, including the Sentient platform. One way a potentially adversarial platform can deviate from the protocol is by falsely claiming the ownership of a model that is not OMLized. For example, this can be achieved by claiming that a response, $M(\tilde{q})$, from a non-OMLized model, $M$, is a fingerprint response for a key, $\tilde{q}$. To prevent this attack, the protocol can require that the fingerprints satisfy some cryptographic relation that cannot be altered after deployment. For example, [91] proposes a novel hash-based approach called Chain & Hash to achieve this goal for fingerprinting LLMs. Such schemes can be seamlessly applied within the current OML 1.0.
@@ B.5 Discussion
B.5.2 Design Space of Fingerprint Functions
For the most common type of paired fingerprints of the form $\{({\rm key},{\rm response})\}$, it is critical that the host does not have access to the fingerprint keys a priori. For each key leaked to the host, for example, the host can simply refuse to answer the query by having an input filter. One fix to this is to increase the number of fingerprints in the model without degrading model utility, which we explored in Fig 11. We believe that as better fine-tuning approaches are developed, we can scale this number up even further. Scaling the fingerprints gives better security as we discuss in Section B.1.3.
@@ B.5.2 Design Space of Fingerprint Functions
Inspired by the literature on model watermarking [54], we propose a scheme to operationalize the above idea. We choose a subset $S_{v}$ of the model vocabulary. We then partition this subset into “red" and “green" words. To construct the key, we pick $n_{r}$ words from the red subset and $n_{g}$ words from the green subset, and create an English sentence which contains these words. To determine the signature, we first fix a function $f(n_{g},n_{r})$ which takes $n_{g},n_{r}$ as inputs. The simplest such function could be $f(x,y)=\mathbb{I}(x>y)$. Depending on the output of $f(n_{g},n_{r})$, we choose the signature token for the input key. Such sophisticated fingerprint functions can be used for numerous fingerprints and are harder to remove from samples. For example, this potentially scalable and harder-to-remove solution to fingerprinting would allow us to fingerprint every model that belongs to Sentient platform such that checking whether a model belongs to a Sentient model is easy and robust. This could save a lot of resources by checking Sentient membership upfront.
B.6 Implementation Details

### v4

@@ C.2 Coalition Attack
An adversarial host who has legitimately acquired multiple OMLized models can launch a notorious attack known as coalition attacks, where multiple OMLized models are used to evade fingerprint detection. One such attack is studied in [18] where common model merging techniques including [103, 37, 107, 110] are used against instructional fingerprinting [105] and watermarking [50]. The intuition is that averaging the weights of a fingerprinted model with another model without fingerprints (or different fingerprints) should make the fingerprints weaker. In the promising preliminary results of [18], the fingerprinting techniques of [105] demonstrated robustness against such attacks; fingerprints persisted through all model merging that preserve utility. However, this is a weak attack and can be significantly strengthened. Note that one implication of this robustness of model merging is that it can be used for trust-free OML as we discuss in Section C.5.1. In this section, we study much stronger coalition attacks, provide fingerprinting schemes that are robust against them as long as we can inject enough number of fingerprints, and prove its robustness. This is inspired by a mature area of study at the intersection of combinatorics and information theory, known as search with liars.
@@
Strong coalition attacks. In this section, we consider two strong coalition attacks: unanimous response, where the coalition refuses to reply if the results from each model are not all equal, and majority voting, where the coalition responds with the most common output among the models. Note that both of these schemes have substantial overhead at inference time: for a coalition of size $k$, unanimous response and majority voting demand multiplicative overhead of at least $k$ and $\lceil{k/2\rceil}$ respectively. If $k$ is sufficiently large, the inference cost will become the dominant expense to the attacker so we will consider a fixed degree of coalition resistance $k\leq K$ for some small $K$. Note that these are stronger coalition attacks than the simple model merging studied in [18], which simply merges the weights of the $k$ models; even when each model has distinct fingerprints, model merging attack has been demonstrated to fail. The standard fingerprint schemes are robust against model merging attacks as we show in Section C.5.1. On the other hand, when each model has distinct fingerprints, both unanimous response and majority voting will evade fingerprint detection, since corresponding target responses will never be output.
@@
To address these stronger coalition attacks of unanimous response and majority voting, we design a novel fingerprinting scheme. This is inspired by the literature on search with liars, and we show that, with enough fingerprints, we can provably identify the models participating in the coalition attacks. The main idea is to add each fingerprint to multiple OMLized models in a carefully designed manner, such that we can iteratively narrow down the candidate set of deployed OMLized models that contains all the models in the coalition of interest. Precisely, let the total number of possible deployed OMLized models be $N$ and the maximum coalition size is $K$ (or 2K12K-1 in the case of majority voting).
Proposition C.1.
@@
There exists a randomized fingerprinting scheme for a universe of $N$ models which can identify a unanimous response coalition of size $K$ (or a majority voting coalition of size 2K12K-1) using
@@
Proof C.2 (Proof of Proposition C.1).
The scheme proceeds with leave-one-out fingerprinting for partitioning of the models as follows: In each round, we assume the candidate models have been split into $K+1$ disjoint partitions $P_{1},\ldots,P_{K+1}$ such that $N]=P_{1}\sqcup\cdots\sqcup P_{K+1}$. Then, for each partition $P_{i}$, we inject one fingerprint $F_{i}$ into each model in the complement $N]\setminus P_{i}$. When testing for the fingerprint, we check for all $K+1$ possible fingerprints $F_{i}$. This guarantees that there will be a fingerprint $F_{i^{*}}$ which spans the coalition (or the acting majority in the case of majority voting), since the no more than $K$ models that determined the coalition’s output can span at most $K$ distinct partitions. Once we have identified $F_{i^{*}}$, we can eliminate the partition $P_{i^{*}}$ from the candidate set. Our goal will be to recursively apply this procedure until the exact coalition has been identified.
@@
[…]
@@
To shrink, $C$ from size $N$ to $N_{0}$, it is sufficient to have $R_{0}=\log(\frac{N}{N_{0}})/\log(1+\frac{1}{2k+1})=O(K\log N)$ successful rounds. By a binomial tail bound, $O(R_{0}\log(1/\delta))$ rounds are sufficient to guarantee $R_{0}$ successes with probability at least $1-\delta/2$. Now, considering the regime where $C$ is shrinking from size $N_{0}$ to 0 (at worst), we note that
@@
In this regime, we define a round a successful if the candidate set shrinks by at least 1. The only way a round can fail is when all partitions that do not contain any coalition members (of which there must be at least one) do not intersect with $C$. From the above, we see that the round must succeed with probability at least $\frac{1}{K+1}$. Now, to successfully identify the coalition, $N_{0}$ successful rounds suffice (we will terminate early once the coalition is identified). By a binomial tail bound, $O(K\cdot N_{0}\log(1/\delta))$ rounds are sufficient to guarantee $N_{0}$ successes with probability at least $1-\delta/2$. Combining the rounds from both regimes, we see that $R=O\big((K\log N+K^{3}\log K)\log(1/\delta)\big)$ ensures overall success with probability at least $1-\delta$. Finally, recall that each round uses $O(K)$ fingerprints.
Worst-case coalition attacks. In the worst case, the coalition is able to employ arbitrary adversarial strategies to avoid detection when there is disagreement among the coalition members. This is significantly more challenging as the adaptive detection algorithm of Proposition C.1 does not guarantee accurate detection anymore. In general, this problem can be formulated as search with lies [48, 73, 49]. In particular, it follows from [49] that there is no fingerprinting procedure that can deterministically guarantee the identification of the coalition, even when assigning unique fingerprints to all possible subsets of models. (Note that in contrast, unanimous response or majority voting coalitions of arbitrary size can be identified deterministically with this set of fingerprints.) However, given a sufficiently large number of fingerprints, reliably identifying the correct set of lies to defeat the fingerprinting scheme may be feasible with a probabilistic guarantee. We demonstrate this in the following proposition.
Proposition C.3.
@@
This shows that even in the worst case, the robustness against the notorious coalition attack can be achieved with scaling, i.e., as long as we have enough fingerprints. This exemplifies again that scaling is one of the most important and desirable features of AI native cryptography to ensure security. Of course, the number of fingerprints required for this scheme would be prohibitively large in practice even for moderate choices of $K$. Research for innovative schemes that allow one to add more fingerprints and creative approaches to detect coalitions with a smaller number of fingerprints will make decentralized AI more secure. At the same time, we believe this result can be improved with a robust version of an adaptive algorithm similar to the one in Proposition C.1. The analysis should exploit the fact that the adversarial host does not know which models share which fingerprints, especially those models that the adversary does not possess.
Proof C.4 (Proof of Proposition C.3).
@@
Noting that $K\leq\sqrt{N/2}$ and choosing $M=O(K\log(N/\delta))$ completes the proof.
C.3 The Auditing Protocol under Decentralized and Untrusted Provers
In OML 1.0, we say a protocol is secure if a host who does not acquire signed permission strings when using an OMLized model can be detected with high probability. Ideally, we want a protocol that is secure without relying on trusted provers. Given a pool of decentralized provers, we demonstrate that the auditing protocol is secure as long as at least one of the provers is honest and the fingerprint responses are kept secret.
Threat model. Consider the scenario of Section C.1.1 where model owners, model hosts, and provers interact using the auditing protocol, with one difference: we have a pool of potentially untrusted provers. Concretely, under the threat model of Section C.1.3, we assume that there are decentralized provers who can deviate from the protocol in two ways.
@@ C.3 The Auditing Protocol under Decentralized and Untrusted Provers
Secondly, an adversarial prover can fabricate a proof of usage to frame an honest host. When an adversarial prover reports a fabricated key-response pair, $)(\tilde{q},M.{\rm oml}(\tilde{q}$)), without querying the host, the previous auditing protocol that trusts provers has no way of telling whether the prover is lying or the host has not acquired the signed permission.
[…]
@@
The first attack by adversarial provers colluding with a host is handled by ($i$) and (iiiiii). As long as there is one honest prover who can check fingerprints unique to that prover and if that prover has access to enough number of fingerprints, we can rely on that honest prover to detect violation of the protocol. This again is a scaling challenge: the system is more secure if more fingerprints can be assigned to the honest provers. As long as we have enough fingerprints assigned to the honest provers, robustness of our fingerprints to input perturbation (Section C.1.3) and fine-tuning (Section C.1.4) will still hold.
[…]
@@
For coalition attacks, our schemes in Section C.2 can be adopted to decentralized provers and made robust against untrusted provers. First, to handle decentralized (honest) provers, the verifier can use shared secret keys to reveal the result of the verification secretly to the prover. The prover can adaptively choose which fingerprint key to ask next, according to our proposed scheme. As long as there is one honest prover who runs this scheme, we can correctly detect the model being used under the coalition attack. Note that an adversarial prover can only cause false negatives, i.e., turn a positive proof of usage into a negative proof. The non-adaptive fingerprinting scheme of Proposition C.3 is naturally robust against false negatives, as long as the honest prover makes enough queries. The adaptive fingerprinting scheme of Proposition C.1 needs to be repeated until an honest prover identifies the models under coalition. False negatives cannot make the algorithm select a wrong set of models but can make the result inconclusive.
C.4 Achieving Loyalty in OML 1.0
@@
The auditing protocol for OML 1.0 introduced in this paper addresses Openness and Monetization, but not Loyalty. One of the most important applications of loyalty is the alignment of LLMs to human safety preferences. Recent advances in hardening the models to be robustly aligned against fine-tuning and jail-breaking attacks can shed light on how to achieve Loyalty on top of OML 1.0.
@@
C.5 Discussion
C.5.1 Trust-free OML 1.0
Ideally, we want OML to not rely on the trust of any party, including the auditing platform. One way a potentially adversarial platform can deviate from the protocol is by falsely claiming the ownership of a model that is not OMLized. For example, this can be achieved by claiming that a response, M(q~)M($\tilde${q}), from a non-OMLized model, $M$, is a fingerprint response for a key, $\tilde{q}$. To prevent this attack, the protocol can require that the fingerprints satisfy some cryptographic relation that cannot be altered after deployment. For example, [85] proposes a novel hash-based approach called Chain & Hash to achieve this goal for fingerprinting LLMs. Such schemes can be seamlessly applied within the current OML 1.0.
@@ C.5 Discussion
C.5.2 Design Space of Fingerprint Functions
For the most common type of paired fingerprints of the form $\{({\rm key},{\rm response})\}$, it is critical that the host does not have access to the fingerprint keys a priori. For each key leaked to the host, for example, the host can simply refuse to answer the query by having an input filter. One fix to this is to increase the number of fingerprints in the model without degrading model utility, which we explored in Fig 10. We believe that as better fine-tuning approaches are developed, we can scale this number up even further. Scaling the fingerprints gives better security as we discuss in Section C.1.3.
@@ C.5.2 Design Space of Fingerprint Functions
Inspired by the literature on model watermarking [50], we propose a scheme to operationalize the above idea. We choose a subset $S_{v}$ of the model vocabulary. We then partition this subset into “red" and “green" words. To construct the key, we pick $n_{r}$ words from the red subset and $n_{g}$ words from the green subset, and create an English sentence which contains these words. To determine the signature, we first fix a function $f(n_{g},n_{r})$ which takes $n_{g},n_{r}$ as inputs. The simplest such function could be $f(x,y)=\mathbb{I}(x>y)$. Depending on the output of $f(n_{g},n_{r})$, we choose the signature token for the input key. Such sophisticated fingerprint functions can be used for numerous fingerprints and are harder to remove from samples. For example, this potentially scalable and harder-to-remove solution to fingerprinting would allow us to fingerprint every model that belongs to auditing platform such that checking whether a model belongs to the platform is easy and robust. This could save a lot of resources by checking the membership upfront.
C.6 Implementation Details